Five fair dice are rolled simultaneously

a.) If we add all the numbers that show up, find the probability that we have at least a sum of seven. (Please show your work for this part :)!)
Answer #1

Five fair dice are rolled

so if we say that their sum is S so S has a minimum value of 5.

so we have too find that

Pr(S >= 7) = 1 - Pr(S = 5) - Pr(S = 6)

Pr(S = 5) = Pr(All dices show up the number 1) = (1/6)5 = 1/7776

P(S= 6) = Pr(Four dices shows number 1 and one shows number 2) = 4C1 (1/6)5 = 1/1944

Pr(S >= 7) = 1 - 5/7776 = 7771/7776

> P(S = 6) should be = 5C1 (1/6)^5 = 5/7776
So, P(S ≥ 7) = 1 - 6/7776 = 1 - 1296 = 1295/1296 (ANSWER)

Tulsiram Garg Wed, Sep 29, 2021 2:12 AM
Answer #2

SOLUTION :


5 dices are rolled simultaneously. 

P( Sum ≥ 7) 

= 1 - P(Sum < 7)


Now possible outcomes for sum < 7 are :


(1,1,1,1,1), (1, 1,1, 1, 2) ; (1,1,1,2,1), (1, 1, 2, 1, 1), (1,2, 1, 1, 1) and (2, 1,1,1,1)


There are 6 outcomes.


P (each outcome) = (1/6)^5 


P (sum < 7) 

= 6 * (1/6)^5

= 1/6^4

= 1/1296


Hence, 


P( sum ≥ 7) 1 - 1/1296 = 1295/1296 = 0.999228395 (ANSWER).


answered by: Tulsiram Garg

> Please correct as follows : P( sum ≥ 7) = 1 - 1/1296 = 1295/1296 = 0.999228395 (ANSWER).

Tulsiram Garg Wed, Sep 29, 2021 1:43 AM